# Nick Stracke    Stefan Andreas Baumann11footnotemark: 1    Kolja Bauer11footnotemark: 1    Frank Fundel    Björn Ommer

### v1

@@ 1 Introduction
To understand why that is the case, we take a look at how diffusion models are trained: a varying amount of noise is added to a clean input image (forward process) and the model is tasked to remove the noise from the image (backward process). The amount of added noise is dependent on the diffusion timestep. As a result, the model learns to operate on noisy images and also becomes dependent on the noise timestep as different noise levels require the model to perform different tasks [4, 1]. Since noisy images inherently contain less information than clean images (c.f. Figure 2), we hypothesize that this harms the internal feature representation of diffusion models  [8] and, thus, the extractable world knowledge. Furthermore, the timestep acts as a hyperparameter that influences the internal feature representation and needs to be picked independently for every downstream application (c.f. Figure 13).
@@
We achieve this by viewing a diffusion model as a family of $T$ feature extractors that operate on images with different noise levels and provide features with different characteristics. We consolidate all $T$ feature extraction functions in our feature extractor by aligning internal representations. Specifically, we initialize our feature extractor as a trainable copy of the diffusion model; fine-tune it with clean images and no timestep as input; and align its features with all $T$ timestep-dependent feature extractors of the diffusion model.
@@ 2 Related Work
Diffusion models [52, 21, 53] are generative models that have defined the state-of-the-art in image generation [11, 35, 44, 46, 39, 13], video generation [12, 40], and audio generation [25, 14] in recent years. Their primary purpose is to generate high-quality samples (images, videos, etc.). However, generation can also be interpreted as a pretext task for learning expressive features, since the model has to build up comprehensive world knowledge in order to generate plausible samples [54]. Features from diffusion models (typically referred to as diffusion features) are obtained by passing a noised image through the diffusion model and extracting intermediate feature representations. They have been shown to be useful for a variety of tasks such as finding semantic correspondences [54, 18, 29], semantic and panoptic segmentation [2, 55, 56, 33], classification [26], and object detection [7].
For semantic correspondence matching, features are leveraged to identify semantically matching regions across images. Existing approaches utilize diffusion features either in a zero-shot setting [54, 18] or fine-tune them for the semantic correspondence task [29, 27, 59]. Some zero-shot approaches do not fine-tune on semantic correspondence but still require tuning a prompt to activate attention maps at the query location of the correspondence [18]. In contrast, our approach aims to provide universal features that can be used for various down-stream tasks in a true zero-shot manner.
Fused Features
It has been shown that diffusion features and features from other self-supervised learning methods such as DINOv2 [37] complement each other [58]. DINOv2 features provide sparse but accurate semantic information, while Diffusion features provide dense spatial information albeit with sometimes inaccurate semantics. State-of-the-art approaches for semantic correspondence detection exploit this and fuse features [58, 59]. Compared to DINOv2 features, diffusion features yield smoother, spatially more coherent correspondences [58].
@@ 3.1 Preliminaries
We hypothesize that diffusion features extracted from noisy images $\mathbf{x}_{t}$ encode the noise $\boldsymbol{\epsilon}$ in addition to information from the image. We investigate this hypothesis in a very simple setting by examining how well features $\mathrm{feat}(\boldsymbol{\epsilon};T)$ extracted from the additive noise $\boldsymbol{\epsilon}$ by itself approximate the features $\mathrm{feat}(\mathbf{x}_{t}=\sqrt{\alpha_{t}}\mathbf{x}_{0}+\sqrt{1-\alpha_{t}}\boldsymbol{\epsilon};t)$. Using least squares, we fit a single scalar approximation coefficient to obtain the optimal reconstruction. We then quantify how much of the variance of the overall features is explained by this approximation (see Figure 3). Even at relatively low timesteps, such as $t=261$ used by DIFT [54], encoding pure noise explains a substantial fraction of the features’ variance. Current diffusion feature methods extract this information jointly with the image information. Our proposed method addresses this issue by eliminating the noise from the feature extraction process.
We further analyze the residual and similarly decompose it via the features predicted at $t=0$ (the clean image). We find that they do not fully explain the remainder of the features either. Instead, a substantial part of the feature variance at medium noise timesteps falls to components not present at $t=0$ or $t=T$. This matches observations by previous works [54] that found diffusion features at higher timesteps offer better semantics, despite the added noise at the image input.
@@
In this section, we present our method to extract noise-free general-purpose diffusion features from a pretrained diffusion backbone through a lightweight fine-tuning process. We call our approach CleanDIFT because it extracts clean DIffusion FeaTures. An overview of our setup is shown in Figure 4.
@@ 4 Experiments
In this section, we investigate the utility of our extracted CleanDIFT features. We test our hypothesis that the proposed extraction setup enables us to leverage more of the world knowledge inherent in diffusion models compared to existing diffusion feature extraction methods while being task-agnostic and timestep-independent. To that end, we evaluate our features on a wide range of downstream tasks: unsupervised zero-shot semantic correspondence, monocular depth estimation, semantic segmentation, and classification. We compare our features against standard diffusion features, methods that combine diffusion features with additional features, and non-diffusion-based approaches.
@@ 4.1 Experimental Setup
Following previous works [54, 58, 59], we evaluate our method on a Stable Diffusion (SD) backbone [44]. We apply our method to SD 1.5 and SD 2.1 to enable fair comparisons with existing methods that use either. We fine-tune our feature extraction model for only 400 steps, taking 30 minutes on a single A100 GPU. We extract features after the U-Net’s middle block and after each of the U-Net’s decoder blocks, except the two final blocks. A detailed visualization of where we extract features is provided in the Appendix. This yields a total of $K=11$ feature maps that we align between the diffusion model and the feature extraction model. Our point-wise feature projection heads consist of three stacked Feed Forward Networks (FFNs). We study the effect of different projection head architectures in the Appendix. We train with a batch size of 8, a learning rate of 2e-6 with a linear warmup, and use Adam [24] as our optimizer. For the stratified timestep sampling, we utilize $I=3$ stratification bins across all our experiments, i.e. three different noise levels per training image.
@@
We fine-tune our feature extraction model on a subset of COYO-700M [5], which is similar to the LAION [48] dataset on which Stable Diffusion 1.5 and 2.1 were trained on originally. That way, we can ensure that all performance improvements originate from the feature extraction model consolidating the diffusion model’s internal feature representations over time and not from choosing a different dataset that matches the test dataset distribution more closely. The subset consists only of images with a minimum size of $512^{2}$. We crop and resize all images to match the corresponding input resolution of the underlying diffusion model.
@@ 4.2 Unsupervised Semantic Correspondence
As many previous diffusion feature methods focus on (unsupervised) semantic correspondence matching [54, 58, 59, 29], we perform an extensive evaluation of our method on this task. Following previous works on semantic correspondence matching [54, 58, 59], we measure our performance in Percentage of Correct Keypoints (PCK). We average PCK directly across all keypoints, not over images. We use $\alpha=0.1$ as a threshold and report both PCK values with error margins relative to the image size and to the bounding box size, denoted as $\text{PCK}_{\text{img}}$ and $\text{PCK}_{\text{bbox}}$ respectively. We evaluate the performance on the test split of the SPair-71k dataset, which consists of approximately 12k image pairs from 18 categories. Some existing works [54, 58, 59] evaluate on additional datasets but find SPair-71k to be the most challenging and therefore the most informative benchmark.
@@ Results
We first compare our extracted features to DIFT [54], an approach that detects semantic correspondences using standard diffusion features. Substituting these with our CleanDIFT features yields a performance increase of $1.79$ percentage points for $\text{PCK}_{\text{img}}$ and $1.86$ percentage points for $\text{PCK}_{\text{bbox}}$. Notably, DIFT averages the extracted feature maps across 8 different noise samples. Without this averaging over noise samples, our performance gain is even larger. This indicates that our feature extraction model learns more than a mere averaging over the noise in the diffusion model’s feature maps. We present an extended version of the time-step dependent performance analysis conducted by [54] in Figure 5: We evaluate the diffusion model’s performance for different timesteps $t$ in two settings. In the first setting, we provide the diffusion model with a noisy input image $\mathbf{x}_{t}$ as usual. In the second setting, we demonstrate that feeding the clean image along with a non-zero timestep is not a viable solution to obtain meaningful features: We provide the diffusion model with the clean input image $\mathbf{x}_{0}$ for all timesteps $t$. We observe that the model’s performance for the clean input image degrades faster and has a lower peak than for the noisy input. This is to be expected, as the diffusion model was trained on noisy images, not clean images. Our CleanDIFT features are timestep-independent and significantly outperform the standard diffusion features, even after tuning for an optimal noising timestep.
A Tale of Two Features [58] extends the approach of DIFT by combining diffusion features with DINOv2 [37] features. Again, we replace the standard diffusion features with our CleanDIFT features and observe that the performance gain transfers when combining our features with DINOv2 features. Telling Left from Right [59] further improves upon the results of A Tale of Two Features by introducing a test-time adaptive pose alignment strategy. We observe that the performance gain transfers to this setting as well. To the best of our knowledge, Telling Left from Right combined with our CleanDIFT features sets a new state-of-the-art in unsupervised zero-shot semantic correspondence matching. In summary, replacing standard diffusion features with our CleanDIFT features consistently results in a significant performance improvement across all three methods. An overview of the results is provided in Tab. 1 and a more extensive evaluation per category is provided in the Appendix.
@@ 4.4 Semantic Segmentation
To further investigate the difference between standard noisy diffusion features and our CleanDIFT features, we evaluate on the semantic segmentation task by training linear probes on our CleanDIFT features and on standard diffusion features. We utilize SD2.1 as the diffusion backbone and extract features at the same location as [54]. This procedure yields feature maps of size $48^{2}$. We train our linear probe on the $48^{2}$ feature maps and upscale the obtained segmentation masks using nearest neighbor upsampling. We train and evaluate on the PASCAL VOC dataset [15]. Following common practice [33], we use mean Intersection over Union (mIOU) as the evaluation metric. Qualitative results are shown in Figure 9. Using our features, we observe significantly less noisy feature maps than with standard diffusion features. A quantitative comparison of our CleanDIFT feature’s performance against standard diffusion features across timesteps is provided in Figure 8. Notably, the optimal timestep appears to be around $t=100$, in contrast to the optimal timestep for semantic correspondences, which [54] found to be $t=261$. This highlights the need for tuning a timestep individually per downstream task. Our method both alleviates the need for such a timestep tuning and outperforms the standard diffusion features for the optimal timestep.
@@
Figure 9: Qualitative results for semantic segmentation from diffusion features on Pascal VOC [15]. Standard SD features use $t=100$ as the timestep, which we found to perform best quantitatively (c.f. Figure 8). Note how the CleanDIFT segmentation maps are far less noisy compared to those of the standard diffusion features.
@@ 4.5 Classification
To assess the impact of our method on non-spatial tasks, we evaluate classification performance using pooled features. Pooling mitigates the influence of localized noise, so we anticipate classification performance to remain on par with standard diffusion features unless our setup introduces detrimental effects. We perform k-Nearest Neighbor (kNN) classification with $k=10$ on ImageNet1k [10], using SD 1.5 as the diffusion backbone. We sweep across feature maps and timesteps $t$ for the base model, with results presented in Figure 10. Our analysis shows that the feature map with the lowest spatial resolution, i.e., feature map #0 (see Figure 11), achieves the highest classification accuracy. Furthermore, the optimal timestep $t$ for the base model varies between feature maps. For the best-performing feature map #0, timestep $t=100$ yields the highest classification accuracy. Importantly, CleanDIFT features slightly outperform the standard diffusion features even when using an optimal timestep $t$ for the base model showing that it does not introduce any detrimental effects.
@@ 4.6 Ablation Studies
For simplicity, we perform our ablation studies using DIFT [54] and evaluate the performance for unsupervised zero-shot semantic correspondence matching on a subset of the SPair71k test split.
@@
Table 3: Ablation Study Results. We evaluate the feature extraction models’ performance for zero-shot semantic correspondence matching on a subset of the SPair71k test split. PCK is aggregated per point.
@@ 5 Conclusion
In this paper, we have proposed CleanDIFT, a novel feature extraction method for diffusion features. CleanDIFT yields noise-free, timestep-independent, general-purpose diffusion features. We achieve this by consolidating the timestep-dependent feature representations of a pre-trained diffusion backbone into a single feature representation. To this end, we align the features of our feature extraction model and the pre-trained diffusion backbone in a lightweight fine-tuning process. The training takes approximately 30 min on a single A100 GPU. Our feature extraction model operates on clean images, thus eliminating the need to destroy information by adding noise to the input image. Additionally, our method alleviates the need to tune a timestep per downstream task. In addition, our model does not require ensembling over noise or timesteps, instead it only requires a single forward pass at inference time to extract features. This reduces the inference cost to a fraction compared to methods that employ ensembling or inversion. We evaluate our CleanDIFT features on a wide variety of downstream tasks and demonstrate significant performance improvements over standard diffusion features.

### v2

@@ 1 Introduction
To understand why that is the case, we take a look at how diffusion models are trained: a varying amount of noise is added to a clean input image (forward process) and the model is tasked to remove the noise from the image (backward process). The amount of added noise is dependent on the diffusion timestep. As a result, the model learns to operate on noisy images and also becomes dependent on the noise timestep as different noise levels require the model to perform different tasks [4, 1]. Since noisy images inherently contain less information than clean images (cf. Figure 2), we hypothesize that this harms the internal feature representation of diffusion models [9] and, thus, the extractable world knowledge. Furthermore, the timestep acts as a hyperparameter that influences the internal feature representation and needs to be picked independently for every downstream application (cf. Figure 14).
@@
We achieve this by viewing a diffusion model as a family of $T$ feature extractors that operate on images with different noise levels and provide features with different characteristics. We consolidate all $T$ feature extraction functions in our feature extractor by aligning their internal representations. Specifically, we initialize our feature extractor as a trainable copy of the diffusion model; fine-tune it with clean images and no timestep input; and align its features with all $T$ time-dependent feature extractors of the diffusion model.
@@ 2 Related Work
Diffusion models [60, 24, 61] are generative models that have defined the state-of-the-art in image generation [12, 41, 51, 53, 46, 14], video generation [13, 47], and audio generation [30, 15] in recent years. Their primary purpose is to generate high-quality samples (images, videos, etc.). However, generation can also be interpreted as a pretext task for learning expressive features, since the model has to build up comprehensive world knowledge in order to generate plausible samples [62, 17, 9, 32, 27]. Features from diffusion models (typically referred to as diffusion features) are obtained by passing a noised image through the diffusion model and extracting intermediate feature representations. They have been shown to be useful for a variety of tasks such as finding semantic correspondences [62, 21, 35], semantic and panoptic segmentation [2, 63, 64, 39], classification [31], and object detection [8].
For semantic correspondence matching, features are leveraged to identify semantically matching regions across images. Existing approaches utilize diffusion features either in a zero-shot setting [62, 21] or fine-tune them for the semantic correspondence task [35, 33, 67]. Some zero-shot approaches do not fine-tune on semantic correspondence but still require tuning a prompt to activate attention maps at the query location of the correspondence [21]. In contrast, our approach aims to provide universal features usable for various down-stream tasks in a true zero-shot manner.
Further, diffusion features have been shown to complement features from other self-supervised learning methods such as DINOv2 [66, 43, 18]. DINOv2 features provide sparse but accurate semantic information, while Diffusion features provide dense spatial information albeit with sometimes inaccurate semantics. State-of-the-art approaches for semantic correspondence detection exploit this and fuse features [66, 67]. Compared to DINOv2 features, diffusion features yield smoother, spatially more coherent correspondences [66].
@@ 3.1 Preliminaries
We hypothesize that diffusion features extracted from noisy images $\mathbf{x}_{t}$ encode the noise $\boldsymbol{\epsilon}$ in addition to information from the image. We investigate this hypothesis in a very simple setting by examining how well features $\mathrm{feat}(\boldsymbol{\epsilon};T)$ extracted from only the noise $\boldsymbol{\epsilon}$ approximate the features $\mathrm{feat}(\mathbf{x}_{t}=\sqrt{\alpha_{t}}\mathbf{x}_{0}+\sqrt{1-\alpha_{t}}\boldsymbol{\epsilon};t)$. Using least squares, we fit a single scalar approximation coefficient to obtain the optimal reconstruction. We then quantify how much of the variance of the overall features is explained by this approximation (cf. Figure 3). Even at relatively low timesteps, such as $t=261$ used by DIFT [62], encoding pure noise explains a substantial fraction of the features’ variance. Current diffusion feature methods extract this information jointly with the image information. Our proposed method addresses this issue by eliminating the noise from the feature extraction process.
We further analyze the residual and similarly decompose it via the features predicted for the clean image $\mathrm{feat}(x_{t}=x_{0};t=0)$. We find that they do not fully explain the remainder of the features either. Instead, a substantial part of the feature variance at medium noise timesteps is timestep dependent and cannot be attributed to components present at $t=0$ or $t=T$. This matches observations by previous works [62] that found diffusion features at higher timesteps offer better semantics, despite the added noise at the image input.
@@
We present CleanDIFT, our method to address the problem of noisy and time-dependent diffusion features. CleanDIFT extracts clean DIffusion FeaTures from a pretrained diffusion backbone through a lightweight fine-tuning process. An overview of our setup is shown in Figure 4.
@@ 4 Experiments
We test our hypothesis that the proposed extraction setup enables us to leverage more of the world knowledge inherent in diffusion models compared to existing diffusion feature extraction methods while being task-agnostic and timestep-independent. To that end, we evaluate our features on a wide range of downstream tasks: unsupervised zero-shot semantic correspondence, monocular depth estimation, semantic segmentation, and classification. We compare our features against standard diffusion features, methods that combine diffusion features with additional features, and non-diffusion-based approaches.
@@ 4.1 Experimental Setup
Following previous works [62, 66, 67], we evaluate our method on a Stable Diffusion (SD) backbone [51]. We apply our method to SD 1.5 and SD 2.1 to enable fair comparisons with existing methods that use either. We fully fine-tune our feature extraction model on image-caption pairs for only 400 steps, taking 30 minutes on a single A100 GPU, which was sufficient for our strong performance and further training did not yield any significant gains. We extract features after the U-Net’s middle block and after each of the U-Net’s decoder blocks, except the two final blocks. A detailed visualization of where we extract features is provided in Figure 11. This yields a total of $K=11$ feature maps that we align between the diffusion model and the feature extraction model. Our point-wise feature projection heads consist of three stacked Feed Forward Networks (FFNs) that are zero-initialized such that initially they act as identity mappings due to their residual connections. Since every aligned feature map has its own projection head, this results in 45M additional trainable parameters for SD 2.1. We study the effect of different projection head architectures in Sec. E. We train using Adam [29] with a batch size of 8 and a learning rate of 2e-6 with a linear warmup. For stratified timestep sampling, we utilize $I=3$ stratification bins across all our experiments, i.e. three different noise levels per training image.
@@
We fine-tune our feature extraction model on a random subset of COYO-700M [5], which is similar to the LAION [56], the dataset that Stable Diffusion 1.5 and 2.1 were trained on originally. That way, we ensure that all performance improvements originate from the feature extraction model consolidating the diffusion model’s internal feature representations over time, not from choosing a different dataset that matches the test dataset distribution more closely. The subset selects images with a minimum size of $512^{2}$. We crop and resize them to match the corresponding input resolution of the underlying diffusion model. We analyze the effect of using different datasets in Sec. G.
@@ 4.2 Unsupervised Semantic Correspondence
As many previous diffusion feature methods focus on (unsupervised) semantic correspondence matching [62, 66, 67, 35], we perform an extensive evaluation of our method on this task. Following previous works on semantic correspondence matching [62, 66, 67], we measure our performance in Percentage of Correct Keypoints (PCK). We average PCK directly across all keypoints, not over images. We use $\alpha=0.1$ as a threshold and report both PCK values with error margins relative to the image size and to the bounding box size, denoted as $\text{PCK}_{\text{img}}$ and $\text{PCK}_{\text{bbox}}$ respectively. We evaluate the performance on the test split of the SPair-71k dataset, which consists of approximately 12k image pairs from 18 categories. Some existing works [62, 66, 67] evaluate on additional datasets but find SPair-71k to be the most challenging and therefore the most informative benchmark. For the text prompt we use “A photo of a category.”, with category being the corresponding category of the SPair image. We experiment with distilling the text conditioning in Sec. B.
@@ Results
We first compare our extracted features to DIFT [62], an approach that detects semantic correspondences using standard diffusion features. Substituting these with our CleanDIFT features yields a performance increase of $1.79$ absolute percentage points for $\text{PCK}_{\text{img}}$ and $1.86$ percentage points for $\text{PCK}_{\text{bbox}}$. Notably, DIFT averages the extracted feature maps across 8 different noise samples. Without this averaging over noise samples, our performance gain is even larger (2.81 PCK@$\alpha_{\textrm{img}}$ gain). This indicates that our feature extraction model learns more than a mere averaging over the noise in the diffusion model’s feature maps (see Figure 5 and Sec. C for examples). We present an extended version of the time-step dependent performance analysis conducted by [62] in Figure 6: We evaluate the diffusion model’s performance for different timesteps $t$ in two settings. In the first setting, we provide the diffusion model with a noisy input image $\mathbf{x}_{t}$ as usual. In the second setting, we demonstrate that feeding the clean image along with a non-zero timestep is not a viable solution to obtain meaningful features: We provide the diffusion model with the clean input image $\mathbf{x}_{0}$ for all timesteps $t$. We observe that the model’s performance for the clean input image degrades faster and has a lower peak than for the noisy input. This is to be expected, as the diffusion model was trained on noisy images, not clean images. Importantly, our CleanDIFT features are timestep-independent and consistently outperform standard diffusion features, even when the latter are optimized for the best-performing timestep. We further observe that this advantage generalizes effectively to other backbones, such as DiTs [44] (see Tab. 7).
A Tale of Two Features [66] extends the approach of DIFT by combining diffusion features with DINOv2 [43] features. Again, we replace the standard diffusion features with our CleanDIFT features and observe that the performance gain transfers when combining our features with DINOv2 features. Telling Left from Right [67] further improves upon the results of A Tale of Two Features by introducing a test-time adaptive pose alignment strategy. We observe that the performance gain transfers to this setting as well. To the best of our knowledge, Telling Left from Right combined with our CleanDIFT features sets a new state-of-the-art in unsupervised zero-shot semantic correspondence matching. In summary, replacing standard diffusion features with our CleanDIFT features consistently results in a significant performance improvement across all three methods. We show an overview of the results in Tab. 1 and a more extensive evaluation per category in Sec. B.
@@ 4.4 Semantic Segmentation
To further investigate the difference between standard noisy diffusion features and our CleanDIFT features, we evaluate on the semantic segmentation task by training linear probes on our CleanDIFT features and on standard diffusion features. We utilize SD2.1 as the diffusion backbone and extract features at the same location as [62]. This procedure yields feature maps of size $48^{2}$. We train our linear probe on the $48^{2}$ feature maps and upscale the obtained segmentation masks using nearest neighbor upsampling. We train and evaluate on the PASCAL VOC dataset [16]. Following common practice [39], we use mean Intersection over Union (mIOU) as the evaluation metric. Qualitative results are shown in Figure 8. Using our features, we observe significantly less noisy segmentations than with standard diffusion features. We show a quantitative comparison of our CleanDIFT feature’s performance against standard diffusion features across timesteps in Figure 9. Notably, the optimal timestep appears to be around $t\!=\!100$, in contrast to the optimal timestep for semantic correspondences, which [62] found to be $t=261$. This highlights the need for tuning a timestep individually per downstream task. Our method both alleviates the need for such a timestep tuning and outperforms the standard diffusion features for the optimal timestep.
@@
Figure 8: Qualitative results for semantic segmentation from diffusion features on Pascal VOC [16]. Standard SD features use $t=100$ as the timestep, which we found to perform best quantitatively (cf. Figure 9). Note how the CleanDIFT segmentation maps are far less noisy than those of the standard diffusion features.
@@ 4.5 Classification
To assess the impact of our method on non-spatial tasks, we evaluate classification performance using pooled features. Pooling mitigates the influence of localized noise, so we anticipate classification performance to remain on par with standard diffusion features unless our setup introduces detrimental effects. We perform k-Nearest Neighbor (kNN) classification with $k=10$ on ImageNet1k [11], using SD 1.5 as the diffusion backbone. We sweep across feature maps and timesteps $t$ for the base model, with results presented in Figure 10. Our analysis shows that the feature map with the lowest spatial resolution, i.e., feature map #0 (see Figure 11), achieves the highest classification accuracy. Furthermore, the optimal timestep $t$ for the base model varies between feature maps. For the best-performing feature map, $t=100$ yields the highest classification accuracy.
@@
Importantly, CleanDIFT features slightly outperform the standard diffusion features even when using an optimal timestep $t$ for the base model showing that it does not introduce any detrimental effects.
@@ 4.6 Ablation Studies
For simplicity, we perform our ablation studies using DIFT [62] and evaluate the performance for unsupervised zero-shot semantic correspondence matching on a subset of the SPair71k [37] test split.
@@
Table 3: Ablation Study Results. We evaluate the feature extraction models’ performance for zero-shot semantic correspondence matching on the SPair71k test split. PCK is aggregated per point.
@@ 5 Conclusion
In this paper, we introduced CleanDIFT, a novel approach for extracting diffusion features. CleanDIFT produces noise-free, timestep-independent, general-purpose diffusion features by consolidating timestep-dependent representations from a pre-trained diffusion backbone into a unified feature representation. We achieve this alignment between our feature extraction model and the pre-trained diffusion backbone through a lightweight fine-tuning procedure that takes approximately 30 minutes on a single A100 GPU. Operating directly on clean images, our method eliminates the information loss associated with adding noise to input images. Furthermore, CleanDIFT removes the requirement for tuning timesteps for each downstream task and avoids the computational overhead of ensembling over noise levels or timesteps. Instead, our method efficiently extracts features with just a single forward pass at inference time, substantially reducing inference costs compared to methods relying on ensembling or inversion. Extensive evaluations of CleanDIFT across diverse downstream tasks demonstrate significant performance improvements over conventional diffusion features.